OpenStudy (anonymous):

What line is perpendicular to 3X+7Y=2?

OpenStudy (anonymous):

which have slope 7/3

OpenStudy (anonymous):

What are the choices?

OpenStudy (anonymous):

A)-7x-3y=2 B)-7x+3y=2 C)3x-7y=2 D)-3x+7y=2 @ilikemath50 @zaibali.qasmi

OpenStudy (anonymous):

its B...

OpenStudy (anonymous):

@zaibali.qasmi thank you so much!!

OpenStudy (anonymous):

its okay...